# RMT-KD: Random Matrix Theoretic Causal Knowledge Distillation

### v3

@@ I Introduction and Prior Works
Recent advances in natural language processing (NLP) and computer vision rely on increasingly large deep learning models such as BERT [1] and convolutional networks such as ResNet [2]. While highly accurate, these models incur substantial inference latency, memory footprint, and carbon emissions [3]. Classical compression methods include knowledge distillation (KD), which transfers predictions from a frozen teacher to a smaller student [4], with refinements such as DistilBERT [5] and progressive shrinking [6]; pruning, which removes unimportant weights or channels [7]; and low-rank factorisation, which compresses convolutional filters [8] and adaptation matrices in large language models [9]. Yet these methods depend on heuristic thresholds, yield hardware-unfriendly sparsity, or lack a principled statistical rule.
In this paper, we ask: Is there a principled statistical basis for model reduction? We argue that Random Matrix Theory (RMT) provides such a foundation [10]. In high-dimensional settings, the eigenvalue spectrum of activation covariances typically separates into a bulk that reflects random noise and a small number of outliers that capture structured, causal features [11, 12]. This separation, formalized by the Marchenko–Pastur (MP) law [13], offers a rigorous criterion for identifying informative directions, as shown in [14].
@@
Figure 1: Architecture of RMT-KD for iterative distillation. At each stage, hidden layer activations are analyzed with RMT principles to identify causal directions, followed by projection and self-distillation. This process is repeated across layers until the benefits saturate.
@@ II RMT-KD: Methodology
We propose an iterative self-distillation method that compresses neural networks by progressively reducing their width under RMT guidance. Unlike standard knowledge distillation, where a student learns from a fixed teacher, our approach uses a single model that distills itself at intermediate checkpoints. Each iteration treats the current model as teacher and its reduced counterpart as student, trained with a combined loss:
@@
Figure 2: The empirical eigenvalue distribution of the covariance matrices computed on the calibration dataset for BERT-base on the SST dataset.
@@
Experimental Setup: We trained all models from scratch: a 12-layer Transformer identical to BERT-base (139M parameters), its 6-layer TinyBERT counterpart (44M), and ResNet-50 (23M). Experiments ran on a single NVIDIA RTX 6000 GPU with CUDA, with GPU power measured using vendor tools. Datasets followed standard splits: GLUE and AG News for language tasks, and CIFAR-10 for vision. Tokenization and preprocessing adhered to best practices.
@@
Figure 3: (a, top) Accuracy vs. parameter reduction and (b, bottom) power consumption vs. inference speedup on GLUE datasets ($\sigma^{2}=$ median eigenvalue, initial quantile = 50%).
@@ III Results and Discussion
Fig. 2 shows the evolution of empirical eigenvalue spectra of covariance matrices computed on calibration data from different depths of BERT-base trained on SST (GLUE). After the validation accuracy threshold is reached, activations encode structured information rather than random noise, and the spectra deviate markedly from the Wigner semicircle law. In the embedding block, most eigenvalues cluster near zero, with only a few outliers exceeding the Marchenko–Pastur bulk (red dashed line), indicating that most directions are noise-dominated while a small subset carries meaningful signal. Deeper layers exhibit broader spectra with smoother decay and a larger fraction of eigenvalues above the bulk edge, reflecting the accumulation and propagation of task-relevant information. This trend is consistent with the idea that later representations are more specialized and semantically structured. Importantly, the bulk cutoff is computed adaptively from the data using RMT rather than a fixed threshold as in PCA, ensuring that only statistically significant causal directions are retained at each step. This adaptive criterion supports our iterative approach: early layers can be compressed aggressively, while deeper layers with richer spectra need conservative reduction, as their distributions deviate increasingly from the Wigner Semicircle Law.
Fig. 3a compares accuracy and parameter counts before and after RMT-based iterative compression. Across all model–dataset pairs, accuracy remains within 2–3 percentage points of the baseline despite substantial reduction, and in cases such as BERT-base on SST, performance even improves slightly—likely due to removal of noisy or redundant parameters. Parameter counts drop sharply, with BERT-base reduced by about 80% while retaining near-original accuracy, highlighting its overparameterization and the ability of RMT-based reduction with self-distillation to preserve informative directions. BERT-tiny still shows 58% reduction, suggesting redundancy even in smaller models, though with less dramatic impact. ResNet-50 exhibits the smallest reduction, around 48%, reflecting its more compact convolutional structure. Overall, the results indicate that larger, more overparameterized models benefit most from statistically guided compression, while smaller architectures are closer to their limits.
Figure 4: Comparison of memory on disk and energy efficiency for all models and GLUE datasets, $\sigma^{2}=MedianEigenvalue$, initial quantile = 50%
@@
Figure 5: Accuracy–reduction tradeoff for BERT-base (GLUE) and ResNet-50 (CIFAR) as a function of the eigenvalue quantile used to initialize $\sigma^{2}$. The x-axis shows the quantile, the y-axis shows accuracy (decreasing) and parameter reduction (increasing). The best balance occurs near 40%.
Ablation Study: Fig. 5 shows the trade-off between accuracy and parameter reduction when varying the quantile for $\sigma^{2}$ initialization in BERT (GLUE) and ResNet (CIFAR-10). At low quantiles, accuracy remains close to the baseline but reductions are limited; at high quantiles, aggressive pruning causes sharp accuracy loss. The best balance occurs around the 40%–50% quantile, near the median, where both architectures retain high accuracy with substantial compression. This regime is particularly effective for reducing inference time and energy while preserving performance. ResNet offers greater reduction potential, while BERT is constrained by fixed embedding and token projection layers, which dominate beyond the 40% quantile and limit compression without accuracy degradation.
@@
Table I: Comparison of KD methods. BERT results are on GLUE; ResNet-50 results are on CIFAR-10. Theseus = BERT-of-Theseus, PKD = Patient Knowledge Distillation, AT = Attention Transfer, FitNet = Hints for Thin Deep Nets, CRD = Contrastive Representation Distillation.
@@ IV Conclusion
We introduce RMT-KD, an iterative distillation framework that combines layer-wise RMT analysis with self-distillation to compress deep models while preserving accuracy. Unlike pruning or heuristic truncation, RMT-KD offers a statistically grounded, data-driven rule for dimensionality reduction that retains dense causal structure. Experiments on BERT (GLUE, AG News) and ResNet (CIFAR-10), it achieves up to 80% parameter reduction with only 2% accuracy loss, 2.8× faster inference, and 80% lower energy use. The models remain hardware-efficient since projections are dense and avoid sparse kernels. Gains are largest for overparameterized transformers, while smaller models like ResNet-50 show modest improvements, reflecting proximity to their efficiency frontier. Performance also depends on calibration subset quality, which may limit robustness under distribution shift.

### v4

@@ I Introduction and Prior Works
Recent advances in natural language processing (NLP) and computer vision rely on increasingly large deep learning models such as BERT [6] and convolutional networks such as ResNet [9]. While highly accurate, these models incur substantial inference latency, memory footprint, and carbon emissions [21]. Classical compression methods include knowledge distillation (KD), which transfers predictions from a frozen teacher to a smaller student [10], with refinements such as DistilBERT [19] and progressive shrinking [4]; pruning, which removes unimportant weights or channels [8]; and low-rank factorisation, which compresses convolutional filters [12] and adaptation matrices in large language models (LLMs) [11]. Yet these techniques often depend on heuristic thresholds, yield hardware-unfriendly sparsity, or lack a principled statistical rule.
In this paper, we ask: Is there a principled statistical basis for model reduction? We argue that Random Matrix Theory (RMT) provides such a foundation [3]. In high-dimensional settings, the eigenvalue spectrum of activation covariances typically separates into a bulk that reflects random noise and a small number of outliers that capture structured, causal features [16, 15]. This separation, formalized by the Marchenko–Pastur (MP) law [14, 7], offers a rigorous criterion for identifying informative directions.
@@
Fig. 1: Architecture of RMT-KD for iterative distillation. At each stage, hidden layer activations are analyzed with RMT principles to identify causal directions, followed by projection and self-distillation. This process is repeated across layers until the benefits saturate.
@@ II RMT-KD: Methodology
We propose an iterative self-distillation method that compresses neural networks by progressively reducing their width under RMT guidance. Unlike standard knowledge distillation, where a student learns from a fixed teacher, our approach uses a single model that distills itself at intermediate checkpoints. Each iteration treats the current model as the teacher and its reduced counterpart as the student, trained with a combined loss:
@@
Experimental Setup: We trained all models from scratch: a 12-layer Transformer identical to BERT-base (139M parameters), its 6-layer TinyBERT counterpart (44M), and ResNet-50 (23M). Experiments ran on a single NVIDIA RTX 6000 GPU with CUDA, with GPU power measured using vendor tools. Datasets followed standard splits, preprocessing and tokenization: GLUE for language tasks, and CIFAR-10 for vision. For BERT, reported results on GLUE refer to the average performance across SST-2, QQP, and QNLI datasets.
Fig. 2: The empirical eigenvalue distribution of the activation matrices computed on the calibration dataset for BERT-base on the SST dataset.
@@
Fig. 3: (a, top) Accuracy vs. parameter reduction and (b, bottom) power consumption vs. inference speedup on GLUE datasets ($\sigma^{2}=$ median eigenvalue, initial quantile = 50%).
@@ III Results and Discussion
Fig. 2 shows the evolution of empirical eigenvalue spectra of activation matrices computed on calibration data from different depths of BERT-base trained on SST (GLUE). After the validation accuracy threshold is reached, activations encode structured information rather than random noise, and the spectra deviate markedly from the Wigner semicircle law. In the embedding block, most eigenvalues cluster near zero, with only a few outliers exceeding the Marchenko–Pastur bulk (red dashed line), indicating that most directions are noise-dominated while a small subset carries meaningful signal. Deeper layers exhibit broader spectra with smoother decay and a larger fraction of eigenvalues above the bulk edge, reflecting the accumulation and propagation of task-relevant information. This trend is consistent with the idea that later representations are more specialized and semantically structured. Importantly, the bulk cutoff is computed adaptively from the data using RMT rather than a fixed threshold as in PCA, ensuring that only statistically significant causal directions are retained at each step. This adaptive criterion supports our iterative approach: early layers can be compressed aggressively, while deeper layers with richer spectra need conservative reduction, as their distributions deviate increasingly from the Wigner Semicircle Law.
Fig. 3a compares accuracy and parameter counts before and after RMT-based iterative compression. Across all model–dataset pairs, accuracy remains within 2–3 percentage points of the baseline despite substantial reduction, and in cases such as BERT-base on SST, performance even improves slightly—likely due to removal of noisy or redundant parameters. Parameter counts drop sharply, with BERT-base reduced by about 80% while retaining near-original accuracy, highlighting its overparameterization and the ability of RMT-based reduction with self-distillation to preserve informative directions. BERT-tiny still shows 58% reduction, suggesting redundancy even in smaller models, though with less dramatic impact. ResNet-50 exhibits the smallest reduction, around 48%, reflecting its more compact convolutional structure. Overall, the results indicate that larger, more overparameterized models benefit most from statistically guided compression, while smaller architectures are closer to their efficiency limits.
Fig. 4: Comparison of memory on disk and energy efficiency for all models and GLUE datasets, $\sigma^{2}=MedianEigenvalue$, initial quantile = 50%
@@
Fig. 5: Accuracy–reduction tradeoff for BERT-base (GLUE) and ResNet-50 (CIFAR) as a function of the eigenvalue quantile used to initialize $\sigma^{2}$. The x-axis shows the quantile, the y-axis shows accuracy (decreasing) and parameter reduction (increasing). The best balance occurs near 40%.
Ablation Study: Fig. 5 shows the trade-off between accuracy and parameter reduction when varying the quantile for $\sigma^{2}$ initialization in BERT (GLUE) and ResNet (CIFAR-10). At low quantiles, accuracy remains close to the baseline but reductions are limited; at high quantiles, aggressive pruning causes sharp accuracy loss. The best balance occurs around the 40%–50% quantile, near the median, where both architectures retain high accuracy with substantial compression. This regime is particularly effective for reducing inference time and energy while preserving performance. ResNet offers greater reduction potential, while BERT is constrained by fixed embedding and token projection layers, which dominate beyond the 40% quantile and limit further compression without accuracy degradation.
@@
Table I: Comparison of KD methods. Theseus = BERT-of-Theseus, PKD = Patient Knowledge Distillation, AT = Attention Transfer, FitNet = Hints for Thin Deep Nets, CRD = Contrastive Representation Distillation.
@@ IV Conclusion
We introduce RMT-KD, an iterative distillation framework that combines layer-wise RMT analysis with self-distillation to compress deep models while preserving accuracy. Unlike pruning or heuristic truncation, RMT-KD offers a statistically grounded, data-driven rule for dimensionality reduction that retains dense causal structure. Experiments on BERT (GLUE) and ResNet (CIFAR-10), it achieves up to 80% parameter reduction with only 2% accuracy loss, 2.8× faster inference, and 80% lower energy use. The models remain hardware-efficient since projections are dense and avoid sparse kernels. Gains are largest for overparameterized transformers, while smaller models like ResNet-50 show modest improvements, reflecting proximity to their efficiency frontier. Performance also depends on calibration subset quality, which may limit robustness under distribution shift.